### v2

# CADDesigner: Conceptual Design of CAD Models Based on General-Purpose Agent

Abstract
Computer-Aided Design (CAD) plays a pivotal role in industrial manufacturing but typically requires a high level of expertise from designers. To lower the entry barrier and improve design efficiency, we present an agent for CAD conceptual design powered by large language models (LLMs). The agent accepts both abstract textual descriptions and freehand sketches as input, engaging in interactive dialogue with users to refine and clarify design requirements through comprehensive requirement analysis. Built upon a novel Context-Independent Imperative Paradigm (CIP), the agent generates high-quality CAD modeling code. During the generation process, the agent incorporates iterative visual feedback to improve model quality. Generated design cases are stored in a structured knowledge base, enabling continuous improvement of the agent’s code generation capabilities. Experimental results demonstrate that our method achieves state-of-the-art performance in CAD code generation.
Figure 1: Demonstration of various CAD models generated by CADDesigner. Our method supports multimodal input and a broad range of CAD operations, including extrusion, revolution, fillet/chamfer, sweeping, lofting, etc., as well as the creation of standard components such as flanges and screws.
Introduction
Computer-Aided Design (CAD) software is widely used across production and manufacturing processes in a range of industries, including automotive, aerospace, and medical device sectors. Its adoption has significantly improved industrial productivity and design efficiency. Traditional CAD platforms—such as OnShape, AutoCAD, SolidWorks, and CATIA—typically rely on manual modeling by skilled designers with deep domain knowledge and technical expertise. However, recent advances in artificial intelligence, particularly the emergence of large language models (LLMs), present promising opportunities to automate CAD model generation, thereby reducing the entry barrier and accelerating the design workflow.
Early research on automatic CAD model generation has primarily focused on parametric modeling approaches (Wu, Xiao, and Zheng 2021; Li et al. 2025c; Khan et al. 2024a, b). However, these methods are constrained by the limited diversity of training data and the representational capacity of their model outputs, supporting only a narrow set of CAD operations. Recent work leveraging large language models (LLMs) for CAD code generation has shown promise in overcoming these limitations (Li et al. 2025a; Wang et al. 2025; Rukhovich et al. 2024). Fine-tuned LLMs can interpret multimodal user inputs and generate corresponding CAD modeling code. Nevertheless, fine-tuning open-source LLMs requires substantial GPU resources, and high-quality CAD training datasets remain scarce, limiting the diversity and quality of the generated code.
To address these challenges, we introduce CADDesigner, an LLM-powered agent for conceptual CAD design. The agent accepts both abstract textual descriptions and hand-drawn sketches as user input, engaging in interactive dialogues with designers to iteratively refine and clarify design requirements through comprehensive requirement analysis.
To enhance the code generation process, we propose a novel Context-Independent Imperative Paradigm (CIP) that enables the agent to produce high-quality CAD modeling code. The agent leverages visual feedback throughout the generation process to incrementally improve model fidelity. Representative results generated by CADDesigner are shown in Figure 1, and our main contributions are summarized as follows:
  • We present CADDesigner, a novel framework for CAD modeling code generation. CADDesigner accepts abstract textual descriptions and sketches as user input, and integrates tools for requirement analysis, knowledge-constrained code generation, and vision-based error correction to generate CAD scripts that satisfy user intent.
  • We introduce a context-independent paradigm for CAD modeling script generation based on an imperative function-calling structure. This paradigm decouples individual CAD operations and improves code generation quality through explicit type annotations, LLM-friendly error representations, and self-evolving capabilities. It supports a wide range of operations, including extrusion, revolution, fillet, chamfer, sweeping, lofting, etc.
The remainder of this paper is organized as follows. We begin with a review of related work. We then introduce our proposed CAD modeling code generation framework, CADDesigner, along with the novel code generation paradigm designed to enhance code quality. Next, we describe our experimental setup and present the results. Finally, we provide a comprehensive analysis and comparison to highlight the strengths and limitations of the proposed approach.
Related Work
This section provides a comprehensive review of three key areas relevant to our work: parametric CAD modeling, LLM-driven CAD generation, and agent-based model generation.
Parametric CAD Model Generation
Parametric CAD model generation approaches commonly employ supervised learning techniques to produce sequences of CAD commands, which can be imported into CAD modeling software to generate editable, parametric files (Willis et al. 2021; Wu, Xiao, and Zheng 2021). Systems such as DeepCAD (Wu, Xiao, and Zheng 2021), Fusion 360 Gallery (Willis et al. 2021), SkexGen (Xu et al. 2022), and Diffusion-CAD (Zhang et al. 2025) adopt customized neural network architectures trained to generate command sequences. However, they primarily support only basic modeling operations such as sketching and extrusion.
Additional efforts (Khan et al. 2024a; Guo et al. 2022; Ma et al. 2024) explore CAD model reconstruction from point cloud data, while more recent methods (Li et al. 2025d; Chen et al. 2025) aim to infer structured, editable CAD models directly from unstructured RGB images (either single- or multi-view), bridging the gap between 2D visual understanding and 3D CAD modeling. Despite these advancements, existing approaches remain limited in their ability to generate complex CAD models, primarily due to constrained training data diversity and limited output representation capacity.
LLM-based CAD Model Generation
Fine-tuned large language models (LLMs) and vision-language models (VLMs) have emerged as a prominent direction for CAD model generation, enabling the synthesis of CAD command sequences or Python code from multimodal input. Cad-LLM (Wu et al. 2023b) and CadVLM (Wu et al. 2024) utilize fine-tuned LLMs to generate parametric models from sketches and images. CAD-MLLM (Xu et al. 2024) targets multimodal inputs—including text, images, and point clouds—for parametric model generation. CAD-Llama (Li et al. 2025a) proposes a hierarchical annotation pipeline that transforms command sequences into semantically structured CAD code. This is followed by adaptive pre-training and instruction tuning, enabling LLMs to generate high-fidelity parametric 3D models. CAD-Recode (Rukhovich et al. 2024) converts point clouds into sequential representations, then uses a pre-trained LLM (Qwen2-1.5B (Team 2024)) to generate corresponding CAD code. Text-to-CadQuery (Xie and Ju 2025) directly maps text descriptions to CadQuery code, leveraging pre-trained LLMs’ capabilities in Python generation and spatial reasoning.
While these methods demonstrate strong generation capabilities, they remain limited by several factors: the computational cost of fine-tuning large models, the closed nature of many commercial LLMs, and the scarcity of high-quality, diverse training datasets for CAD.
Agent-based CAD Model Generation
Another research direction focuses on constructing CAD generation agents using prompt engineering and ultra-large foundation models, circumventing the need for additional training. CAD-Assistant (Mallis et al. 2024) presents a general-purpose CAD agent framework that integrates visual and language models (VLLMs) as planners to generate Python code for FreeCAD, achieving zero-shot capability across diverse CAD tasks. SeekCAD (Li et al. 2025b) combines retrieval-augmented generation (RAG), visual feedback, and a chain-of-thought (CoT) mechanism to generate customized CAD modeling code with self-optimization. 3D-PreMise (Yuan et al. 2024) investigates the potential and limitations of LLMs in program synthesis for manipulating 3D software and generating parametrically controlled shapes. CADCodeVerify (Wu et al. 2023a) incorporates iterative validation and refinement loops, using visual language models to pose and answer verification questions about the generated CAD outputs.
In contrast to these works, CADDesigner accepts multimodal user inputs—text and sketches—to capture detailed design requirements, engages in interactive refinement with users, and enhances code quality and model fidelity through visual feedback and iterative optimization.
CAD Modeling Architecture
Figure 2: The CADDesigner agent follows a ReAct-style paradigm to progressively transform user requirements into valid CAD models through iterative reasoning, tool execution, and feedback refinement. It first refines user requirements into structured specifications, generates executable modeling code using domain APIs, and analyzes execution results via both symbolic (e.g., shell logs, geometric metadata) and visual feedback (e.g., rendered 3D views). In the figure, arrows of different colors indicate information flows across various stages of the modeling loop, while the numbered circles represent the temporal order of data generation. Readers are encouraged to follow the numbers sequentially to understand the agent’s internal workflow.
In the following subsections, we first present the architecture of the CADDesigner agent. We then provide an overview of the CAD tools used, followed by detailed descriptions of each individual tool integrated into the system.
ReAct Agent rather than Workflow
CADDesigner employs a ReAct-style (Yao et al. 2023) agent-centric architecture, where a central agent governs the entire modeling loop through iterative reasoning, tool execution, and feedback integration.
Unlike traditional hard-coded workflows, our architecture leverages prompt-based workflow specification, where system prompts encode procedural guidance for the agent. This design allows for flexible and revisable orchestration, as the workflow is not statically programmed but interpreted at runtime by the agent. As a result, users can intervene at any stage of the process, like adding constraints, correcting directions, or adjusting goals, leading to a truly interruptible and human-in-the-loop modeling cycle
The ReAct-style loop includes reasoning, acting, and reflecting: the agent expands coarse inputs into parameterized specs, generates executable code, renders outputs, evaluates results, and iteratively refines the model until user intent is satisfied. By consolidating control within the agent, we enable adaptive tool use and interpretable, interactive CAD automation in under-specified or evolving design tasks.
CAD Tools
CADDesigner employs a set $\mathcal{T}=\{T_{1},T_{2},T_{3},T_{4}\}$ of four CAD-specific tools, each serving a distinct functional role in the design-to-modeling pipeline:
  • $T_{1}:\mathcal{I}\to\mathcal{D}$ maps user-provided multimodal inputs $\mathcal{I}$ to detailed design descriptions $\mathcal{D}$, refining initial requirements into structured design specifications.
  • $T_{2}:\mathcal{D}\to\mathcal{C}$ translates finalized design descriptions into executable CAD modeling code $\mathcal{C}$. A new code paradigm for CAD code generation has been designed and used by $T_{2}$. This will be discussed in detail in the following section “CAD Code Generation Paradigm”.
  • $T_{3}:\mathcal{C}\to\mathcal{M}$ executes the generated code $\mathcal{C}$ to produce corresponding CAD models $\mathcal{M}$, aligning with standard shell command execution processes.
  • $T_{4}:\{\mathcal{M},\mathcal{D}\}\to(P,F)$ returns a binary tuple where $P\in\{0,1\}$ indicates the termination flag ($P=1$ triggers termination), and $F$ is diagnostic feedback for potential refinement. Inside this tool, an multi-view render result of model $\mathcal{M}$ will be generated. We denote this render result as $\mathcal{V}$. Thus $T_{4}$ can be separated into 2 steps: $T_{4}^{1}:\mathcal{M}\to\mathcal{V}$, and $T_{4}^{2}:\{\mathcal{V},\mathcal{D}\}\to(P,F)$
The sequential composition of these tools forms a functional pipeline:
$T_{4}^{1}\circ T_{3}\circ T_{2}\circ T_{1}:\mathcal{I}\to\mathcal{V}$
This pipeline is augmented with a closed-loop correction mechanism: after generating visual feedback $\mathcal{V}$, the system invokes $T_{4}^{2}$ to obtain $(P,F)$. If $P=0$, the feedback $F$ is used to guide revised design generation via $T_{3}\circ T_{2}\circ T_{1}$. If $P=1$, the process terminates successfully. Thus, the loop continues until $T_{4}^{2}$ returns a termination signal.
Requirements Analysis
During the conceptual design phase, designers typically have only rough descriptions or sketches, making it challenging to precisely define detailed model parameters as well as modeling process. To address this, we support designers by conducting requirement analysis to clarify and refine their design intention.
Let the set of initial information input by the user be $\mathcal{I}(Text,Img)$, where TextText represents simple text description and ImgImg represents some sketchs. Let $\mathcal{P}$ be the set of necessary parameters required for generating a CAD model. The process of the CADDesigner conducting detailed design can be expressed as $F:\mathcal{I}\to\mathcal{D}_{\text{detail}}$, where $\mathcal{D}_{\text{detail}}$ is detailed designs containing specific parameters, satisfying $\mathcal{P}\subseteq\mathcal{D}_{\text{detail}}$, which is accomplished by $T_{1}$. The user’s revision process for the detailed design is an iterative function:
$R:\mathcal{D}_{\text{detail}}\times\mathcal{U}\to\mathcal{D}_{\text{detail}}$
where $\mathcal{U}$ is the set of user revision operations. After $n$ iterations of revision, we obtain $\mathcal{D}_{\text{final}}=R^{n}(\mathcal{D}_{\text{detail}},\mathcal{U})$, which is then passed to the next stage.
Knowledge Constrained Code Generation
$\mathcal{D}_{\text{final}}$ is the detailed design confirmed by the user. The process by which the code generation tool $T_{2}$ receives $\mathcal{D}_{\text{final}}$ and generates CAD modeling code is expressed as $G:\mathcal{D}_{\text{final}}\to\mathcal{C}$. A knowledge base $K$ is constructed, which includes a set of function annotations $Anno\subseteq\mathcal{K}$ and a set of basic model cases $Case\subseteq\mathcal{K}$. The process by which the CADDesigner generates code with reference to the knowledge base can be expressed as
$G^{\prime}:\mathcal{D}_{\text{final}}\times\mathcal{K}\to\mathcal{C^{\prime}}$
and $\mathcal{C^{\prime}}\supset\mathcal{C}$, meaning that referring to the knowledge base can improve the success rate of code generation. The running result of the code is delivered to the judgment function $J:\mathcal{C}\to\{0,1\}$. When $J(\mathcal{C})=1$, the code can run successfully and generate a model $\mathcal{M}=T_{3}(\mathcal{C})$, which is then passed to the next stage; when $J(\mathcal{C})=0$, new code is regenerated based on the error feedback.
Visual Based Iterative Error Correction
Multi-view snapshots are taken of the generated CAD model $\mathcal{M}$, resulting in an image set $\mathcal{V}=\{Img_{1},Img_{2},\dots,Img_{6}\}$ (corresponding to the front top left, back bottom right, back top left, front bottom right, top and right views respectively). CADDesigner uses the function
$Check:\mathcal{D}\times\mathcal{V}\to\{0,1\}$
to check whether the model meets the user’s intent. When $Check(\mathcal{D},\mathcal{V})=0$, it returns visual feedback $F$ to the code generation stage to regenerate the code; when $Check(\mathcal{D},\mathcal{V})=1$, the result is delivered to the user for final judgment. The user’s satisfaction with the modeling result is determined by $Sat:\mathcal{M}\to\{0,1\}$. If $Sat(\mathcal{M})=1$, $\mathcal{M}$ is added to the case library $\mathcal{K}$, that is, $Case^{\prime}=Case\cup\{\mathcal{M}\}$; if $Sat(\mathcal{M})=0$, it re-enters the code generation stage. In the entire process, except for the user feedback links ($R$ and SatSat), the state transition function $Trans:State\to State^{\prime}$ (where the state transitions and judgments of StateState are independently completed by the CADDesigner).
CAD Code Generation Paradigm
In this section, we provide a detailed description of a new CAD code generation paradigm from four key aspects: the LLM-friendly Context-Independent Imperative Paradigm (CIP), structured tracking and error information, and the automatic construction of a reusable modeling knowledge base.
Definition
We define two mappings: one between the language system and the semantic space, and the other between two language systems.
Definition 1 (Language System, Express Mapping and Explain Mapping):
Formally, language system is defined as a tuple $\mathcal{L}=(\Sigma^{*},\mathbf{R})$.
where $\Sigma^{*}$ denotes the set of all finite strings over a finite alphabet. $\mathcal{M}$ is the semantic space.
  • Express Mapping $\mathbf{R}$ that converts strings to semantic elements is defined as:
    $\mathbf{R}:\Sigma^{*}\rightarrow\mathcal{M}$
    And,
    $\forall s\in\Sigma^{*},\quad\exists\mathcal{M}^{\prime}\subsetneq\mathcal{M}\quad\text{s.t.}\quad\mathbf{R}(s)\subseteq\mathcal{M}^{\prime}$
  • Explain Mapping $\mathbf{R}^{-1}$ that recovers strings from semantic elements is defined as:
    $\mathbf{R}^{-1}:\mathcal{M}\rightharpoonup\Sigma^{*}$
    This mapping is partial, meaning it is not necessarily defined for all elements in $\mathcal{M}$. Specifically,
    $\exists m\in\mathcal{M}\quad\text{s.t.}\quad\mathbf{R}^{-1}(m)=\varnothing,$
    indicating that certain semantics cannot be explained or reconstructed using strings in the given language system.
Definition 2 (Translation Mapping):
Let $\mathcal{L}_{A}=(\Sigma_{A}^{*},\mathbf{R}_{A})$ and $\mathcal{L}_{B}=(\Sigma_{B}^{*},\mathbf{R}_{B})$ be two language systems over the same semantic space $\mathcal{M}$. The translation mapping from $\Sigma_{A}^{*}$ to $\Sigma_{B}^{*}$ is defined as the composition:
$\mathbf{T}_{A\rightarrow B}:=\mathbf{R}_{B}^{-1}\circ\mathbf{R}_{A},$
where for $s_{A}\in\Sigma_{A}^{*}$, the translation $s_{B}=\mathbf{T}_{A\rightarrow B}(s_{A})$ is the result of first mapping $s_{A}$ into its semantic meaning in $\mathcal{M}$ via $\mathbf{R}_{A}$, then generating a corresponding string $s_{B}\in\Sigma_{B}^{*}$ using $\mathbf{R}_{B}^{-1}$.
Context-Independent Imperative Paradigm
To address the limitations of chaining modeling language such as CadQuery, we propose a new modeling paradigm termed CIP (Context-Independent Imperative Paradigm). We give the following definitions based on the definition above.
Definition 3 (CAD Operation Semantic Space):
The CAD operation semantic space is defined as a set of tuples representing morphisms:
$\forall m$ $\in\mathcal{M}_{\text{CADOP}}$
$m$ $:=(\mathcal{F}_{\textbf{OP}},[\textbf{Obj}],\textbf{Params})$
Here $\mathcal{F}_{\textbf{OP}}$ is a morphism defined as:
$\mathcal{F}_{\textbf{OP}}(\mathcal{O})=\mathcal{O}^{\prime},\quad\text{where }\mathcal{O},\mathcal{O}^{\prime}\in\textbf{Obj}\cup\textbf{Params}$
Obj denotes instances of geometric entities such as Solid, Face, etc., and Params refers to parameters required for a specific CAD operation.
Definition 4 (Context-Independent Imperative Paradigm (CIP)):
We define CIP as a language system:
$\mathcal{L}_{\textbf{CIP}}$ $=(\Sigma^{*}_{\textbf{CIP}},\mathbf{R}_{\textbf{CIP}})$
where:
  • $\forall s\in\Sigma^{*}_{\textbf{CIP}},\quad s:=\texttt{Operation}([\texttt{Obj}],\texttt{Params}),$ $s$ is a string or called a ’statement’,
  • $\forall s\in\Sigma^{*}_{\textbf{CIP}},\ \exists!m\in\mathcal{M}_{\text{CADOP}}\quad\text{s.t.}\quad\mathbf{R}_{\textbf{CIP}}(s):=m=(\mathcal{F}_{\mathbf{OP}},[\textbf{Obj}],\textbf{Params})$
E.g., $s=\texttt{extrude\_rsolid(S, direction, distance)}$, in which extrude_rsolid is an Operation, S is an Object and $\texttt{direction},\texttt{distance}$ are Parameters.
Each statement in CIP forms a self-contained and semantically closed unit. The mapping $\mathbf{R}_{\textbf{CIP}}$ yields a unique operational triple without requiring any execution history or context. This ensures that CIP is a truly context-independent paradigm.
LLM Friendly Error Information
CadQuery’s error messages are typically raw Python exceptions with unstructured stack traces. Due to its reliance on method chaining and implicit states, errors often occur far from their actual cause, providing little semantic guidance. As a result, LLMs struggle to identify how to fix the underlying issue.
In CIP, we introduce fine-grained error handling nearly at the level of each atomic operation. Each failure triggers a structured diagnostic message that includes likely causes and suggested fixes which are derived from patterns observed in common failure cases. This design ensures that errors carry actionable signals, making it much easier for LLMs to locate and repair faults.
Formally, each error is represented as:
$\mathit{ErrMsg}=(\mathit{ErrCau},\ \mathit{ErrLoc},\ \mathit{CorrAct}),$
where $\mathit{ErrCau}$ denotes the root cause of failure, $\mathit{ErrLoc}$ indicates the precise location in the user script, and $\mathit{CorrAct}$ suggests one or more potential corrective actions.
This format enables LLMs to directly extract relevant debugging signals. In addition, corrective hints are programmatically injected based on the error type and recent execution trace, offering high-level semantic suggestions rather than raw exceptions. Since all runtime logs and structured diagnostics are centrally managed, the LLM agent can access and reason over them during iterative code generation.
Explicit Type Annotation
A common source of failure observed in PythonOCC and CadQuery stems from LLMs misinterpreting the return types of certain methods. This is particularly problematic in distinguishing semantically similar but hierarchically distinct geometric entities, such as Wire and Edge. While both may appear interchangeable in conventional geometry contexts, they represent different levels of abstraction in CAD systems, which is a distinction critical for correct modeling behavior.
To address this, CIP emphasizes type clarity through both API naming conventions and type-annotated documentation. Specifically, each API function in CIP follows the pattern:
ActionName_rReturnType
Here, ActionName denotes the CAD operation, r is a fixed prefix indicating “returns”, and ReturnType explicitly declares the type of the returned object. This ensures that even without inspecting documentation, the function signature itself conveys type information, reducing ambiguity during LLM inference.
In addition, each function is accompanied by detailed DocStrings with explicit type annotations, which not only guide human developers but also enhance LLM performance in code-generation tasks by reinforcing accurate type semantics during prompt-based retrieval or few-shot learning.
These design choices are later validated through ablation studies, demonstrating their effectiveness in reducing type-related errors and improving modeling success rates.
Self-Evolving by Composition Operation
Our system supports a comprehensive set of atomic CAD operations, which includes extrusion, filleting, chamfering, sweeping, lofting, etc., covering nearly the entire design operation space. Beyond this foundation, we allow the explicit definition and preservation of composite operations formed by sequences of atomic ones during the modeling process. These composites can be treated as new atomic operations in subsequent tasks, effectively turning hard-to-discover modeling strategies into reusable knowledge components and expanding the agent’s action space with higher-level, user-defined primitives. Base on Definition 4, we can easily get the Corollary that composition operation is also context-independent and bijective. We observed that the agent could directly invoke preserved composite operations to generate parts such as screws and flanges, which significantly improved the success rate of one-shot model generation.
Code Framwork Knowedge Base Construction
To enhance modeling accuracy and reduce repetitive agent failures, we construct a structured knowledge base $\mathcal{K}$ that serves as external memory during code generation. The knowledge base consists of two key components:
  • Function Annotations ($\mathit{Anno}$): semantic descriptions of commands, including parameter formats, return types, and usage notes;
  • Case Examples ($\mathit{Case}$): verified code snippets preserved as reference patterns.
To automate the construction of the knowledge base $\mathcal{K}$, we design a rule-based pipeline that extracts and organizes information directly from the source code. Specifically, we collect function signatures, parameter types, and return types from Python annotations and type hints. These are then aligned with the corresponding natural language descriptions found in the function-level docstrings.
When building the vectorized representation of $\mathcal{K}$, we employ a customized chunking strategy that respects the structural boundaries of the Markdown documents. This design ensures that retrieval at inference time yields coherent and contextually complete knowledge snippets for guiding code generation.
Experiments and Comparison
We evaluate our CAD modeling code generation algorithm on a workstation equipped with an workstation with an 8-core Intel CPU, using Python 3.12 as the runtime environment. The agent system was implemented on a customed framework tailored to our task, and retrieval-augmented knowledge base access is provided by RAGFlow.
As for LLM API service, we use Claude-4-Sonnet as main agent, Gemini-2.5-Pro for code generator, and Gemini-2.5-Flash for refining requirements.
Evaluation Metrics
To evaluate the geometric fidelity of the generated CAD models, we adopt several standard shape similarity metrics, including IoU (Intersection over Union), CD (Chamfer Distance), and HD (Hausdorff Distance). These metrics assess how closely the generated shapes match the ground-truth geometry from different perspectives.
In addition, we introduce three process-oriented metrics to analyze modeling stability and reliability under ablation settings. Pass@1 denotes the proportion of cases where valid code is produced on the first attempt, reflecting generation precision. AVG Re (Average Retry) captures the average number of retries required during the iterative correction loop, indicating the ease of convergence. SUC refers to the percentage of successful cases.
Let $\mathcal{S}$ denote the reference models and $\mathcal{G}$ represent the generated models, with $\mathcal{X}$ and $\mathcal{Y}$ representing their corresponding point clouds, whose mathematical formulations are described as follows:
Intersection over Union (IoU) is utilized to measure the similarity between generated models and the ground truth.
$\mathrm{IoU}(\mathcal{G},\mathcal{S})=\frac{\mathcal{G}\cap\mathcal{S}}{\mathcal{G}\cup\mathcal{S}},$
where $\mathcal{G}\cap\mathcal{S}$ denotes the overlapping volume between the reference and generated models, and $\mathcal{G}\cup\mathcal{S}$ represents their combined volumetric union. A value of 11 indicates perfect alignment, while 0 signifies no overlap.
Chamfer Distance (CD) calculates point-wise proximity between point clouds sampled from $\mathcal{X}$ and $\mathcal{Y}$:
$\mathrm{CD}(\mathcal{X},\mathcal{Y})=\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\min_{y\in\mathcal{Y}}\|x-y\|_{2}^{2}+\frac{1}{|\mathcal{Y}|}\sum_{y\in\mathcal{Y}}\min_{x\in\mathcal{X}}\|y-x\|_{2}^{2}.$
Hausdorff distance(HD) is a metric for quantifying the similarity between two point sets, primarily used for evaluating the resemblance of object contours.
$\mathrm{HD}(\mathcal{X},\mathcal{Y})=\max\left\{\sup_{x\in\mathcal{X}}\inf_{y\in\mathcal{Y}}d(x,y),\sup_{y\in\mathcal{Y}}\inf_{x\in\mathcal{X}}d(x,y)\right\}$
where, $d(x,y)$ is the Euclidean distance between points $x$ and $y$, $\sup$ (supremum) and $\inf$ (infimum) represent the least upper bound and greatest lower bound, respectively.
Ablation Study on CIP Paradigm
To illustrate the benefits of certain design choices in CIP, such as the error handling mechanism that is friendly to large language models (LLMs) and the overall architecture of the CIP system itself, we conducted an ablation study on Text2CAD dataset. Specifically, we evaluated four variants on the same dataset: CadQuery, CIP, CIP without the detailed error system, and CIP without the return type annotations. The results are presented in Table 1.
Table 1: Ablation study of CIP components and comparative evaluation with CadQuery.
Method Pass@1↑ AVG Re↓ SUC↑
CIP 0.44 1.88 100%
CIP w/o Err 0.44 2.67 77.8%
CIP w/o Type 0.33 2.33 88.9%
CadQuery 0.18 4.00 72.7%
The results show that CIP achieves the best overall performance, clearly outperforming CadQuery in all metrics. When the error handling mechanism is removed (CIP w/o Err), the Pass@1 remains at 0.44, but the AVG Re increases from 1.88 to 2.67, and SUC drops to 77.8%. This suggests that while the agent can still generate correct code initially, the lack of detailed feedback slows down convergence during iterative correction.
In contrast, removing type annotations (CIP w/o Type) causes Pass@1 to drop significantly to 0.33, indicating that the initial code generation is more prone to semantic errors without explicit type guidance. Interestingly, AVG Re (2.33) is lower than in the error-free setting but still higher than full CIP, and SUC also drops to 88.9%. This shows that type annotations help encode critical modeling intent into the prompt, but some type errors still get corrected through retries.
Overall, we observe a clear division of roles: error handling helps the agent quickly recover from mistakes during iterations, and type annotations improve first-attempt accuracy by reducing semantic ambiguity.
Comparison Methods
Figure 3: Comparison of generation results across Text2CAD, CADCodeVerify, cadrille, and our method. CADDesigner achieves the best input alignment.
We compared our method with two learning based text-to-CAD generation methods: Text2CAD (Khan et al. 2024b),cadrille (Kolodiazhnyi et al. 2025) and one agent-based method: CADCodeVerify (Wu et al. 2023a).
  • Text2CAD proposes an end-to-end transformer-based auto-regressive network to generate parametric CAD models from input texts.
  • cadrille proposes a two-stage CAD reconstruction pipeline: supervised fine-tuning followed by reinforcement learning.
  • CADCodeVerify proposes an agent that utilizes validation questions and visual feedback for design verification.
We sampled $1k$ instances in an even distribution of model complexity from the $8k$ test set of Text2CAD for comparison. For each instance, we used the abstract text description as input prompts. We then performed inference using their released weights and our method to generate results for metric calculation. As presented in Table 2 and Figure 3, the experimental results show that the model generated by our method achieves the optimal performance. Our method demonstrates the best prompt-result alignment, followed by CADCodeVerify and Text2CAD. Cadrille fails to generate valid outputs due to poor generalization, as it was trained on expert-level descriptions and cannot adapt to abstract-level inputs during inference.
Table 2: Performance comparison of different Text-to-CAD methods under abstract text inputs.
Method IoU↑ CD↓ HD↓ SUC↑
Text2CAD 0.1831 0.1475 0.5680 96.6%
cadrille 0.0274 0.2162 0.5817 98.2%
CADCodeVerify 0.2348 0.2329 0.4892 86.1%
CADDesigner 0.2769 0.1097 0.4347 100.0%
Limitations
Although our method demonstrates considerable flexibility, it still faces challenges in generating models that require precise geometric constraints. While input text or images can explicitly specify numerical parameters (e.g., dimensions), current large language models (LLMs) and vision-language models (VLMs) exhibit limited understanding and reasoning abilities of spatial geometric relationships and topological dependencies. This often results in outputs that satisfy local constraints (e.g., correct height or diameter of a cup) but violate global consistency (e.g., cup’s handle is inside the cup).
Due to limitations in the current code-generation models, the approach underperforms in generating CAD models involving complex mathematical computations, such as involute gears. These tasks, requiring calculations of fundamental geometric parameters, meshing relationships, or strength validation, yet cannot perform out well due to the complexity of generating correct spatial geometry calculation code.
Conclusion and Future Work
We propose a novel framework CADDesigner for CAD modeling code generation combined with a context-independent CAD modeling script generation paradigm to generate high-quality CAD modeling scripts. Experiments show that this method can achieve SOTA results in the field of CAD conceptual design.
CADDesigner is highly effective for rapidly prototyping CAD models, capable of generating innovative CAD designs and producing high-quality, complex datasets that align text descriptions with corresponding CAD modeling code. Future work could integrate learning-based methods to accept point clouds and B-rep data as inputs, enabling the model to learn geometric and topological constraints through training. Additionally, incorporating models specialized in spatial geometry reasoning could assist CADDesigner in computing complex CAD parameters, such as those required for gear and crankshaft design.