### v4

Abstract
Computer-Aided Design (CAD) plays a pivotal role in industrial manufacturing but typically requires a high level of expertise from designers. To lower the entry barrier and improve design efficiency, we present an agent for CAD conceptual design powered by large language models (LLMs). The agent accepts both textual descriptions and sketches as input, engaging in interactive dialogue with users to refine and clarify design requirements through comprehensive requirement analysis. Built upon a novel Explicit Context Imperative Paradigm (ECIP), the agent generates high-quality CAD modeling code. During the generation process, the agent incorporates iterative visual feedback to improve model quality. Generated design cases are stored in a structured knowledge base, enabling continuous improvement of the agent’s code generation capabilities. Experimental results demonstrate that our method achieves state-of-the-art performance in CAD code generation.
keywords:
[…]
Figure 1: Demonstration of various CAD models generated by CADDesigner. Our method supports multimodal input and a broad range of CAD operations, including extrusion, revolution, fillet/chamfer, sweeping, lofting, etc., as well as the creation of standard components such as flanges and screws.
1 Introduction
Computer-Aided Design (CAD) software is widely used across production and manufacturing processes in a range of industries, including automotive, aerospace, and medical device sectors. Its adoption has significantly improved industrial productivity and design efficiency. Traditional CAD platforms—such as OnShape, AutoCAD, SolidWorks, and CATIA—typically rely on manual modeling by skilled designers with deep domain knowledge and technical expertise. However, recent advances in artificial intelligence, particularly the emergence of LLMs, present promising opportunities to automate CAD model generation, thereby reducing the entry barrier and accelerating the design workflow.
Early research on automatic CAD model generation has primarily focused on parametric modeling approaches [wu2021deepcad, li2025mamba, khan2024cad, text2cad]. However, these methods are constrained by the limited diversity of training data and the representational capacity of their model outputs, supporting only a narrow set of CAD operations. Recent work leveraging LLMs for CAD code generation has shown promise in overcoming these limitations [li2025cad, wang2025cad, rukhovich2024cad]. Fine-tuned LLMs can interpret multimodal user inputs and generate corresponding CAD modeling code. Nevertheless, fine-tuning open-source LLMs requires substantial GPU resources, and high-quality CAD training datasets remain scarce, limiting the diversity and quality of the generated code.
To address these challenges, we introduce CADDesigner, an LLM-powered agent for conceptual CAD design. The agent accepts both textual descriptions and sketches as user input, engaging in interactive dialogues with designers to iteratively refine and clarify design requirements through comprehensive requirement analysis.
To enhance the code generation process, we propose a novel Explicit Context Imperative Paradigm (ECIP) that enables the agent to produce high-quality CAD modeling code. The agent leverages visual feedback throughout the generation process to incrementally improve model fidelity. Representative results generated by CADDesigner are shown in Figure 1, and our main contributions are summarized as follows:
  • We present CADDesigner, a novel framework for CAD modeling code generation. CADDesigner accepts textual descriptions and sketches as user input, and integrates tools for requirement analysis, knowledge-constrained code generation, and vision-based error correction to generate CAD scripts that satisfy user intent.
  • We introduce an explicit context imperative paradigm for CAD modeling script generation based on an imperative function-calling structure. This paradigm decouples individual CAD operations and improves code generation quality through explicit type annotations, LLM-friendly error representations, and self-evolving capabilities. It supports a wide range of operations, including extrusion, revolution, fillet, chamfer, sweeping, lofting, etc.
The remainder of this paper is organized as follows. We begin with a review of related work. We then introduce our proposed CAD modeling code generation framework, CADDesigner, along with the novel code generation paradigm designed to enhance code quality. Next, we describe our experimental setup and present the results. Finally, we provide a comprehensive analysis and comparison to highlight the strengths and limitations of the proposed approach.
2 Related Work
This section provides a comprehensive review of three key areas relevant to our work: parametric CAD modeling, LLM-driven CAD generation, and agent-based model generation.
2.1 Parametric CAD Model Generation
Parametric CAD model generation approaches commonly employ supervised learning techniques to produce sequences of CAD commands, which can be imported into CAD modeling software to generate editable, parametric files [willis2020fusion, wu2021deepcad]. Systems such as DeepCAD [wu2021deepcad], Fusion 360 Gallery [willis2020fusion], SkexGen [xu2022skexgen], and Diffusion-CAD [zhang2025diffusion] adopt customized neural network architectures trained to generate command sequences. However, they primarily support only basic modeling operations such as sketching and extrusion.
Additional efforts [khan2024cad, guo2022complexgen, ma2024draw] explore CAD model reconstruction from point cloud data, while more recent methods [li2025caddreamer, chen2025cadcrafter] aim to infer structured, editable CAD models directly from unstructured RGB images (either single- or multi-view), bridging the gap between 2D visual understanding and 3D CAD modeling. Despite these advancements, existing approaches remain limited in their ability to generate complex CAD models, primarily due to constrained training data diversity and limited output representation capacity.
2.2 LLM-based CAD Model Generation
Fine-tuned LLMs and vision-language models (VLMs) have emerged as a prominent direction for CAD model generation, enabling the synthesis of CAD command sequences or Python code from multimodal input. Cad-LLM [wu2023cad] and CadVLM [wu2024cadvlm] utilize fine-tuned LLMs to generate parametric models from sketches and images. CAD-MLLM [xu2024cad] targets multimodal inputs—including text, images, and point clouds—for parametric model generation. CAD-Llama [li2025cad] proposes a hierarchical annotation pipeline that transforms command sequences into semantically structured CAD code. This is followed by adaptive pre-training and instruction tuning, enabling LLMs to generate high-fidelity parametric 3D models. CAD-Recode [rukhovich2024cad] converts point clouds into sequential representations, then uses a pre-trained LLM (Qwen2-1.5B [team2024qwen2]) to generate corresponding CAD code. Text-to-CadQuery [xie2025text] directly maps text descriptions to CadQuery code, leveraging pre-trained LLMs’ capabilities in Python generation and spatial reasoning.
While these methods demonstrate strong generation capabilities, they remain limited by several factors: the computational cost of fine-tuning large models, the closed nature of many commercial LLMs, and the scarcity of high-quality, diverse training datasets for CAD.
2.3 Agent-based CAD Model Generation
Another research direction focuses on constructing CAD generation agents using prompt engineering and ultra-large foundation models, circumventing the need for additional training. CAD-Assistant [mallis2024cad] presents a general-purpose CAD agent framework that integrates visual and language models (VLLMs) as planners to generate Python code for FreeCAD, achieving zero-shot capability across diverse CAD tasks. SeekCAD [li2025seek] combines retrieval-augmented generation (RAG), visual feedback, and a chain-of-thought (CoT) mechanism to generate customized CAD modeling code with self-optimization. 3D-PreMise [yuan20243d] investigates the potential and limitations of LLMs in program synthesis for manipulating 3D software and generating parametrically controlled shapes. CADCodeVerify [wu2023visual] incorporates iterative validation and refinement loops, using visual language models to pose and answer verification questions about the generated CAD outputs.
In contrast to these works, CADDesigner accepts multimodal user inputs—text and sketches—to capture detailed design requirements, engages in interactive refinement with users, and enhances code quality and model fidelity through visual feedback and iterative optimization.
3 CAD Modeling Architecture
Figure 2: The Intelligent CAD Orchestrator Agent, CADDesigner, follows a ReAct-style paradigm to progressively transform user requirements into valid CAD models through iterative reasoning, tool execution, and feedback refinement. It first refines user requirements into detailed designs, generates executable modeling code using domain APIs, and analyzes execution results via both symbolic (e.g., shell logs and errors) and visual feedback (e.g., rendered 3D views).
In this section, we introduce the CADDesigner architecture, which features a ReAct-style agent interacting with a diverse set of integrated CAD tools to iteratively generate and refine CAD models. The overall framework is illustrated in Figure 2.
3.1 ReAct Agent rather than Workflow
CADDesigner employs a ReAct-style [yao2023react] agent-centric architecture, where a central agent governs the entire modeling loop through iterative reasoning, tool execution, and feedback integration.
Unlike traditional hard-coded workflows, our architecture leverages prompt-based workflow specifications, where the system prompts encode procedural guidance for the agent. This design allows for flexible and revisable task orchestration, as the workflow is not statically programmed but interpreted at runtime by the agent. As a result, users can intervene at any stage of the process, like adding constraints, correcting directions, or adjusting goals, leading to a truly interruptible and human-in-the-loop modeling cycle.
The ReAct-style loop includes reasoning, acting, and reflecting: the agent expands coarse inputs into parameterized specs, generates executable code, renders outputs, evaluates results, and iteratively refines the model until user intent is satisfied. By consolidating control within the agent, we enable adaptive tool use and interpretable, interactive CAD automation in under-specified or evolving design tasks.
3.2 CAD Tools
The intelligent orchestration relies on a diverse suite of CAD-oriented tools. Each tool addresses a specific stage within the design-to-modeling pipeline and works in synergy under the direction of the ReAct agent. This subsection describes the CAD tools integrated in our pipeline.
3.2.1 Core CAD Toolset
CADDesigner employs various CAD-specific tools to support its design-to-modeling pipeline. Central to this set are four tools, denoted $\mathcal{T}=\{T_{1},T_{2},T_{3},T_{4}\}$, each fulfilling a distinct role.
  • Requirement Refinement Tool $T_{1}\colon\mathcal{I}\to\mathcal{D}$ maps user-provided multimodal inputs $\mathcal{I}$ to detailed design descriptions $\mathcal{D}$, refining initial requirements into structured design specifications.
  • Code Generation Tool $T_{2}\colon\mathcal{D}\to\mathcal{C}$ translates finalized design descriptions into executable CAD modeling code $\mathcal{C}$. A new code paradigm for CAD code generation has been designed and used by $T_{2}$. This will be discussed in detail in the section 4.
  • Command Execution Tool $T_{3}\colon\mathcal{C}\to\mathcal{M}$ executes the generated Python code $\mathcal{C}$ to produce the corresponding CAD model $\mathcal{M}$ by running shell commands.
  • Visual Feedback Tool $T_{4}\colon\{\mathcal{M},\mathcal{D}\}\to(P,F)$ returns a binary pair where $P\in\{0,1\}$ indicates the termination flag ($P=1$ triggers termination), and $F$ is diagnostic feedback for potential refinement. Inside this tool, a multi-view render result of model $\mathcal{M}$ will be generated. We denote this render result as $\mathcal{V}$. Thus $T_{4}$ can be separated into 2 steps: $T_{4}^{1}:\mathcal{M}\to\mathcal{V}$, and $T_{4}^{2}:\{\mathcal{V},\mathcal{D}\}\to(P,F)$
The sequential composition of these core tools forms the main functional pipeline:
$T_{4}^{1}\circ T_{3}\circ T_{2}\circ T_{1}:\mathcal{I}\to\mathcal{V}.$ (1)
This pipeline is augmented with a closed-loop correction mechanism: after generating visual feedback $\mathcal{V}$, the system invokes $T_{4}^{2}$ to obtain $(P,F)$. If $P=0$, the feedback $F$ is used to guide revised design generation via $T_{3}\circ T_{2}\circ T_{1}$. If $P=1$, the process terminates successfully. Otherwise, the loop continues until $T_{4}^{2}$ returns a termination signal.
3.2.2 Auxiliary Toolset
In addition to the core CAD toolset, CADDesigner integrates several auxiliary tools that support efficient context management and system interaction. These include:
  • SketchPad Tool: A flexible key-value storage tool designed to mitigate the context explosion problem in LLM-based agents. SketchPad allows the agent to store arbitrary intermediate data at any time during the design process, including image paths, reference code snippets, execution results, and more. By storing context data as keyed entries instead of embedding entire data repeatedly, it significantly reduces LLM context size. SketchPad also supports automatic summarization, tag-based search, and automatic expiration of outdated entries, enabling flexible and scalable context management.
  • File Operation Tools: These tools provide fundamental file-reading and file-writing capabilities, allowing the agent to store or load files as required throughout the CAD design pipeline. Similar to SketchPad, file content or references are stored and accessed efficiently to avoid large context overheads.
Together, these auxiliary tools complement the core CAD toolset by enabling robust, efficient, and scalable operation of the CADDesigner agent in complex multi-turn interactions.
3.3 Requirements Analysis
During the conceptual design phase, designers typically have only rough descriptions or sketches, making it challenging to precisely define detailed model parameters as well as the modeling process. To address this, we support designers by conducting requirement analysis to clarify and refine their design intention.
Let the set of initial information input by the user be $\mathcal{I}\subseteq\mathcal{I}_{\mathrm{txt}}\times\mathcal{I}_{\mathrm{img}}$, where $\mathcal{I}_{\mathrm{txt}}$ represents the set of possible text inputs (including empty text $\emptyset$) and $\mathcal{I}_{\mathrm{img}}$ represents the set of possible image inputs (including a null image $\emptyset$). Let $\mathcal{P}$ be the set of necessary parameters required for generating a CAD model. The process of the CADDesigner conducting detailed design can be expressed as $\mathcal{F}_{\text{design}}:\mathcal{I}\to\mathcal{D}_{\text{detail}}$, where $\mathcal{D}_{\text{detail}}$ is detailed designs containing specific parameters, satisfying $\mathcal{P}\subseteq\mathcal{D}_{\text{detail}}$, which is accomplished by $T_{1}$. The user’s revision process for the detailed design is an iterative function:
$R:\mathcal{D}_{\text{detail}}\times\mathcal{U}\to\mathcal{D}_{\text{detail}},$ (2)
where $\mathcal{U}$ is the set of user revision operations. After $n$ iterations of revision, we obtain $\mathcal{D}_{\text{final}}=R^{n}(\mathcal{D}_{\text{detail}},\mathcal{U})$, which is then passed to the next stage.
3.4 Knowledge Constrained Code Generation
$\mathcal{D}_{\text{final}}$ is the detailed design confirmed by the user. The process by which the code generation tool $T_{2}$ receives $\mathcal{D}_{\text{final}}$ and generates CAD modeling code is expressed as $G:\mathcal{D}_{\text{final}}\to\mathcal{C}$. A knowledge base $\mathcal{K}$ is constructed to support code generation. $\mathcal{K}$ includes a set of function annotations $Anno\subseteq\mathcal{K}$ and a set of basic model cases $Case\subseteq\mathcal{K}$, which provide semantic and structural references for the code generation tool. With the assistance of the knowledge base, the code generation process is extended to:
$G^{\prime}:\mathcal{D}_{\text{final}}\times\mathcal{K}\to\mathcal{C}^{\prime},$ (3)
where $\mathcal{C}^{\prime}$ denotes the code generated referencing the knowledge base, which is semantically and structurally richer compared to the code $\mathcal{C}$ generated without using the knowledge base. The execution result of the code is delivered to the judgment function $J:\mathcal{C}^{\prime}\to\{0,1\}$. When $J(\mathcal{C}^{\prime})=1$, the code can run successfully and generate a model $\mathcal{M}=T_{3}(\mathcal{C}^{\prime})$, which is then passed to the next stage; when $J(\mathcal{C}^{\prime})=0$, the error information is fed back along with the knowledge base $\mathcal{K}$ to guide the next iteration of code generation $G^{\prime}$ to produce a new code. This feedback loop helps to iteratively correct the code and improve the success rate. Therefore, the knowledge base provides essential semantic and structural support for code generation, and the iterative process combined with execution feedback further enhances the likelihood of successful code generation.
3.5 Visual-based Iterative Error Correction
Multi-view snapshots of the generated CAD model $\mathcal{M}$ are taken by $T_{4}^{1}$, producing an image set $\mathcal{V}=\{\text{Img}_{1},\allowbreak\text{Img}_{2},\allowbreak\dots,\allowbreak\text{Img}_{6}\}$ (corresponding to the Front-Top-Left, Back-Bottom-Right, Back-Top-Left, Front-Bottom-Right, Top, and Right views respectively). CADDesigner evaluates whether the model meets the user’s intent $\mathcal{D}_{\text{final}}$ using a binary variable $P\in\{0,1\}$. Internally, $P$ is determined by $T_{4}^{2}$, which involves two components:
  • Visual Question Generation $F_{\mathrm{vq}}:\mathcal{V}\times\mathcal{D}_{\text{final}}\to\mathcal{Q}$, which produces a set of targeted visual questions $\mathcal{Q}=\{q_{1},q_{2},\ldots,q_{n}\}$ based on the user’s requirements and the multi-view images $\mathcal{V}$.
  • Visual Feedback Generation $F_{\mathrm{vf}}:\mathcal{V}\times\mathcal{Q}\to(P,F)$, which analyzes the multi-view images $\mathcal{V}$ to answer the questions in $\mathcal{Q}$, and outputs explicit visual feedback $F$ along with the binary decision $P$. When $P=0$, it returns visual feedback $F$ to the code generation stage to regenerate the code; when $P=1$, the result is delivered to the user for final judgment. The user’s satisfaction with the modeling result is determined by $Sat:\mathcal{M}\to\{0,1\}$. If $Sat(\mathcal{M})=1$, $\mathcal{M}$ is added to the case library $\mathcal{K}$, that is, $Case^{\prime}=Case\cup\{\mathcal{M}\}$; if $Sat(\mathcal{M})=0$, it re-enters the code generation stage.
In the entire process, except for the user feedback links ($R$ and $Sat$), the state transition function $Trans:State\to State^{\prime}$ (where the state transitions and judgments of $State$ are independently completed by the CADDesigner).
4 CAD Code Generation Paradigm
In this section, we provide a detailed description of a new CAD code generation paradigm from three key aspects: the LLM-friendly Explicit Context Imperative Paradigm (ECIP), LLM-friendly CAD code design, and the automatic construction of a reusable modeling knowledge base.
4.1 Explicit Context Imperative Paradigm
In recent years, many works employ CadQuery (CQ) code as an intermediate representation for CAD model generation. It adopts a Fluent API design, where operations are expressed via chained method calls to enable a highly readable and compact modeling workflow. This style relies on an implicit internal context $C$ that tracks the current modeling state and is automatically passed between operations. Formally, each operation $f_{k}$ updates the hidden state $C_{k}$ with parameters $p_{k}$:
$C_{k+1}=f_{k}(C_{k},p_{k}),$ (4)
and a chained user expression corresponds to function composition:
$C_{n}=f_{n}\circ f_{n-1}\circ\cdots\circ f_{1}(C_{0}).$ (5)
While concise, this implicit state obscures semantic dependencies, complicating understanding and automated code generation—especially for Large Language Models that struggle to track hidden contexts. In contrast, low-level APIs like PythonOCC require explicit inputs for each operation, yielding clear semantics but verbose, fine-grained code that lacks high-level abstraction and usability.
To address these limitations, we introduce the Explicit Context Imperative Paradigm (ECIP), which enforces explicit passing of all input objects and parameters to each operation. Formally, each operation $g_{k}$ maps explicit input state $S_{k}$ and parameters $p_{k}$ to a new state $S_{k+1}$:
$S_{k+1}=g_{k}(S_{k},p_{k}).$ (6)
Thus, the modeling workflow is represented as a sequence of explicit state transformations:
$S_{1}\!=\!g_{1}(S_{0},p_{1}),\,S_{2}\!=\!g_{2}(S_{1},p_{2}),\,\ldots,\,S_{n}\!=\!g_{n}(S_{n-1},p_{n}),$ (7)
where all intermediate states $S_{k}$ are explicit variables.
Figure 3: Code comparison between CadQuery (left) and ECIP (right). ECIP explicitly passes context and supports standard Python constructs, improving code clarity and flexibility.
Figure 3 illustrates a typical code comparison between CadQuery and ECIP. CadQuery (left) uses chained method calls that implicitly pass context, making the code compact but less flexible for inserting standard Python statements or loops. In contrast, ECIP (right) explicitly passes all inputs, allowing the use of standard Python control flow like $\mathtt{for}$ loops and variable assignments, which enhances modularity, readability, and ease of debugging. Thus, ECIP combines the clarity of low-level APIs with the expressiveness of high-level modeling, greatly improving the accuracy and interpretability of agent-driven CAD code generation.
4.2 LLM-Friendly CAD Code Design
In addition to the core Explicit Context Imperative Paradigm (ECIP) described in Section 4.1, we incorporate several complementary design features aimed at improving Large Language Model (LLM) agents’ ability to accurately generate and debug CAD modeling code.
4.2.1 Structured Error Information
The error handling mechanism of CadQuery typically produces generic or unstructured error messages, which significantly impede automated debugging and code correction. To address this limitation, ECIP introduces fine-grained structured error messages at the atomic operation level, represented as triples:
$\mathtt{ErrMsg}=(\mathtt{ErrCau},\ \mathtt{ErrLoc},\ \mathtt{CorrAct}),$ (8)
where $\mathtt{ErrCau}$ identifies the root cause of the failure, $\mathtt{ErrLoc}$ specifies the exact location of the error in the script, and $\mathtt{CorrAct}$ suggests potential corrective actions. This structured format provides actionable feedback that LLMs can leverage to improve fault localization and enable automated code repair.
4.2.2 Explicit Type Annotations
Type ambiguity commonly leads to erroneous code generation, particularly when dealing with closely related geometric entities such as Wire and Edge. ECIP mitigates this issue by enforcing explicit type annotations through a disciplined naming convention for functions, following the pattern:
$\mathtt{ActionName\_rReturnType},$ (9)
where $\mathtt{ActionName}$ denotes the CAD operation, $\mathtt{r}$ is a fixed prefix indicating "returns", and $\mathtt{ReturnType}$ explicitly declares the type of the returned object. This ensures that even without inspecting documentation, the function signature itself conveys type information, reducing ambiguity during LLM inference. Each function is also accompanied by comprehensive, type-annotated documentation. This explicit type information reduces ambiguity and improves code generation accuracy by LLMs.
4.2.3 Self-Evolving Composite Operations
To facilitate knowledge accumulation and abstraction, ECIP supports the definition of composite operations—higher-level constructs formed by encapsulating sequences of atomic operations. These composites are saved as new operations within the language, enabling agents to invoke them directly in subsequent tasks.
This self-evolving mechanism effectively expands the language’s set of available operations, allowing reuse of complex modeling patterns such as screws, flanges, or other parametric features. Empirically, this contributes to improved one-shot generation success rates by allowing LLM agents to leverage accumulated domain knowledge efficiently.
4.3 Knowledge Base for Code Framework
LLMs possess strong generative and reasoning abilities but often struggle with the fine-grained semantic requirements of CAD APIs, such as precise parameter formats, geometric constraints, and operation-specific return types. These limitations frequently lead to misinterpreted command signatures or hallucinated inputs, resulting in unstable code generation.
To provide explicit, domain-grounded guidance during modeling, we introduce a structured knowledge base $\mathcal{K}$ coupled with retrieval-augmented generation (RAG). Retrieved function annotations and example cases supply accurate API semantics that complement the LLM’s implicit priors, substantially improving generation reliability. The knowledge base $\mathcal{K}$ consists of two key components:
  • Function Annotations ($\mathtt{Anno}$): semantic descriptions of API functions, including parameter formats, return types, and usage notes;
  • Case Examples ($\mathtt{Case}$): Initially composed of manually selected examples, and gradually expanded during use.
To automate the construction of the knowledge base $\mathcal{K}$, we design a rule-based pipeline that extracts and organizes information directly from the source code. Specifically, we collect function signatures, parameter types, and return types from Python annotations and type hints. These are then aligned with the corresponding natural language descriptions provided in the function-level docstrings.
When building the vectorized representation of $\mathcal{K}$, we employ a customized chunking strategy that respects the structural boundaries of the Markdown documents. This design ensures that retrieval at inference time yields coherent and contextually complete knowledge snippets for guiding code generation.
5 Experiments and Comparison
This section details the implementation of CADDesigner and the comprehensive experimental evaluation conducted to assess its performance in CAD modeling code generation.
5.1 Implementation Details
We evaluate our CAD modeling code generation algorithm on a workstation with an 8-core Intel CPU under Python 3.12. The system is implemented on a customized agent framework that leverages retrieval-augmented generation (RAG) via RAGFlow. For embedding-based retrieval, we utilize the embedding model bge-large-zh-v1.5 to encode knowledge base documents. The retrieval employs a hybrid similarity scoring method, where vector similarity and keyword-based similarity are combined with weights of 0.6 and 0.4, respectively. The top 3 most relevant documents (top-k=3) are retrieved for downstream generation. For LLM services, the primary agent and the code generation tool ($T_{2}$) employ Claude-4-Sonnet, while requirement refinement ($T_{1}$) and both visual feedback sub-tools (visual question generation and visual feedback generation) are based on Gemini-2.5-Flash. Inference uses a temperature of 0.7 and nucleus sampling (top-p) of 0.9 to balance diversity and coherence. Detailed prompt designs, semantic chunking strategies, and other related details are described in the Appendix.
5.2 Dataset
Our experiments utilize the large-scale DeepCAD dataset [wu2021deepcad], which contains approximately 178K parametric CAD models based on the sketch-and-extrude representation. Following the approach in SkexGen [xu2022skexgen], we first remove duplicate models to reduce redundancy. Subsequently, we apply a filtering strategy inspired by HNC-CAD [xu2023hierarchical], retaining only those CAD models where the number of solids does not exceed 5, the total number of loops in all sketches is at most 20, the total number of curves in all loops is at most 60, and the total number of commands in the sketch-and-extrude sequence does not exceed 200. This two-stage filtering result in a refined dataset more suitable for downstream modeling tasks. From the filtered training set, we randomly select 1K models to initialize the RAG knowledge base. Deduplication ensures these models do not overlap with those in the test set. Additionally, a conversion script has been developed to transform the parametric command sequences of each model into ECIP code snippets, which are included in the RAG knowledge base as $\mathtt{Case}$. For evaluation, we randomly sample 200 and 1K models from the filtered test set for ablation studies and comparison experiments, respectively. For each sampled model, a single representative image is rendered from a fixed camera viewpoint to provide visual input in the ablation experiments.
5.3 Evaluation Metrics
To assess the geometric fidelity of the generated CAD models, we adopt several standard shape similarity metrics: Intersection over Union (IoU), Chamfer Distance (CD), and Hausdorff Distance (HD). These metrics evaluate how closely the generated geometry aligns with the ground-truth models from multiple perspectives. In addition, we introduce three process-oriented metrics to analyze modeling stability and reliability under ablation settings: Pass@1 denotes the percentage of cases where valid code is generated on the first attempt, reflecting generation accuracy. AVG Re (Average Retry) captures the mean number of retries required during the iterative correction loop, indicating convergence efficiency. SUC denotes the proportion of cases where a syntactically valid CAD model is successfully generated, i.e., models that can be executed and rendered without errors. Note that a syntactically valid model is not necessarily semantically aligned with the input description. Detailed mathematical definitions and metric computation protocols are provided in Appendix.
5.4 Comprehensive Ablation Study
To comprehensively verify the design choices and the effectiveness of different components in our CADDesigner agent, we conducted two groups of ablation studies on the DeepCAD dataset: one focusing on the ECIP paradigm, and another evaluating the impact of CAD tools as well as input modalities.
5.4.1 Effect of ECIP Paradigm
To illustrate the benefits of certain design choices in ECIP, such as the error handling mechanism that is friendly to LLMs and the overall architecture of the ECIP system itself, we conducted an ablation study on a subset of 200 models randomly sampled from the filtered test set described in Section 5.2. Specifically, we evaluated four variants on the same dataset: CadQuery, ECIP, ECIP without the detailed error system, and ECIP without the return type annotations. The results are presented in Table 1.
Table 1: Ablation study of ECIP components and comparative evaluation with CadQuery on 200 test models.
Pass@1↑ AVG Re↓ SUC↑
ECIP 0.46 1.85 100%
ECIP w/o Err 0.42 2.60 79.5%
ECIP w/o Type 0.31 2.30 90.5%
CadQuery 0.17 3.95 71.5%
The results show that ECIP achieves the best overall performance, clearly outperforming CadQuery in all metrics. When the error handling mechanism is removed (ECIP w/o Err), the Pass@1 decreases slightly from 0.46 to 0.42, while the AVG Re increases from 1.85 to 2.60, and SUC drops to $79.5\%$. This suggests that although the agent can still generate correct code initially, the lack of detailed feedback slows down convergence during iterative correction and reduces the overall success rate. In contrast, removing type annotations (ECIP w/o Type) causes Pass@1 to drop significantly from 0.46 to 0.31, indicating that the initial code generation is more prone to semantic errors without explicit type guidance. The AVG Re (2.30) is higher than full ECIP, but still lower than ECIP without error handling, and SUC also drops to $90.5\%$. This shows that type annotations help encode critical modeling intent into the prompt, yet some type errors can still be corrected through retries. Overall, we observe a clear division of roles: error handling helps the agent quickly recover from mistakes during iterations, and type annotations improve first-attempt accuracy by reducing semantic ambiguity.
5.4.2 Effect of CAD Tools and Input Modalities
To further analyze the individual contributions of the key components in CADDesigner, we conducted an additional ablation study. We evaluated seven variants, labeled A to G, which respectively correspond to: A) removal of function annotations $\mathtt{Anno}$ from the knowledge base, B) removal of basic model cases $\mathtt{Case}$, C) disabling the requirement refinement tool $T_{1}$, D) disabling the visual feedback tool $T_{4}$, E) the full CADDesigner pipeline with text-only input, F) full pipeline with image-only input, and G) full pipeline with combined text and image input. Except for variants F and G, all others use text input exclusively.
Table 2: Ablation study of CADDesigner components and input modalities on 200 test models.
IoU↑ CD↓ HD↓
A (w/o $\mathtt{Anno}$)
B (w/o $\mathtt{Case}$) 0.2643 0.1320 0.4918
C (w/o $T_{1}$) 0.2713 0.1344 0.4764
D (w/o $T_{4}$) 0.2871 0.1160 0.4142
E (Text only) 0.3292 0.1148 0.3758
F (Image only) 0.3720 0.1028 0.3606
G (Text + Image) 0.3992 0.0826 0.2606
Table 2 summarizes the quantitative results across multiple evaluation metrics. In Model A, removing semantic API annotations ($\mathtt{Anno}$) prevents the agent from reliably identifying available operations and their usage, leading to complete failure in generating valid models. This indicates that structured function-level annotations play a critical role in grounding agent-based CAD code generation and guiding convergence toward correct solutions.
Models B, C, and D disable key system components—case examples ($\mathtt{Case}$), requirement refinement ($T_{1}$), and visual feedback ($T_{4}$), respectively. Compared to the text-only variant (Model E), all three models exhibit consistent performance degradation across metrics (IoU: 0.2643, 0.2713, 0.2871 vs. 0.3292), suggesting that each component contributes positively to robust and accurate model generation. Specifically, removing structured case examples and requirement refinement increases semantic ambiguity during early-stage generation, while disabling visual feedback primarily impairs iterative correction, as reflected by higher CD and HD values.
Switching the input modality from text-only (Model E) to image-only (Model F) yields a notable performance improvement, suggesting that abstract textual descriptions alone are often insufficient to fully specify geometric details, whereas images provide richer spatial cues for CAD modeling. Model G, which combines both text and image as input, achieves the best results across all metrics, illustrating that textual information can effectively complement visual context by resolving ambiguities difficult to infer from images alone.
Overall, these results indicate that structured API knowledge, coordinated tool usage, and multimodal input jointly contribute to high-fidelity and robust CAD model generation.
5.5 Comparison Methods
We compared our method with two learning-based text-to-CAD generation methods: Text2CAD [text2cad], cadrille [kolodiazhnyi2025cadrille] and one agent-based method: CADCodeVerify [wu2023visual].
  • Text2CAD proposes an end-to-end transformer-based auto-regressive network to generate parametric CAD models from input texts.
  • cadrille proposes a two-stage CAD reconstruction pipeline: supervised fine-tuning (SFT) on large-scale procedurally generated data, followed by reinforcement learning fine-tuning using online feedback.
  • CADCodeVerify proposes an agent that utilizes validation questions and visual feedback for design verification.
For evaluation, we use 1k samples from the filtered test set described in Section 5.2. For each instance, we used the abstract text description as input prompts. We then performed inference using their released weights and our method to generate results for metric calculation. As presented in Table 3 and Figure 4, the experimental results show that the model generated by our method achieves the optimal performance. Our method demonstrates the best prompt-result alignment, followed by CADCodeVerify and Text2CAD. While cadrille achieves a high SUC score, indicating it can produce syntactically valid CAD models in most cases, the visualized results reveal these outputs are typically meaningless with respect to the input instructions (see Figure 4). This failure is due to cadrille’s poor generalization from expert-level to abstract instructions: it generates structurally valid but semantically irrelevant models. In contrast, the other methods generate outputs that are both syntactically and semantically valid to varying degrees, with our method being the most consistent.
Table 3: Performance comparison of different Text-to-CAD methods under abstract text inputs on 1K test models.
Method IoU↑ CD↓ HD↓ SUC↑
Text2CAD 0.1831 0.1475 0.5680 96.6%
cadrille 0.0274 0.2162 0.5817 98.2%
CADCodeVerify 0.2348 0.2329 0.4892 86.1%
CADDesigner 0.2769 0.1097 0.4347 100.0%
Figure 4: Comparison of generation results across Text2CAD, CADCodeVerify, cadrille, and our method. CADDesigner achieves the best input alignment. Text2CAD and CADCodeVerify show moderate performance, while cadrille generates syntactically valid but semantically incorrect outputs due to poor generalization from expert-level training to abstract inputs.
Figure 5: Comparison of CADDesigner and cadrille on image-based inputs. CADDesigner (blue) generates structurally accurate CAD models that closely match the input images by supporting expressive operations such as revolve and high-level patterns, whereas cadrille (green), relying on low-level extrusion, often produces syntactically valid but geometrically less faithful results.
Additionally, we compare our method with cadrille on image-based inputs to further illustrate their respective modeling capabilities. As shown in Figure 5, CADDesigner consistently generates CAD models that closely match the input images, benefiting from a more expressive CAD operation space. In particular, several models require rotationally symmetric structures that are naturally constructed via revolve operations, which are explicitly supported by CADDesigner but not by cadrille, leading to missing or distorted geometries in the latter. Moreover, models with rich geometric details and repetitive structures are more effectively specified using high-level CAD operations such as arrays or patterned constructs. CADDesigner can leverage these higher-level APIs to compactly and precisely generate such structures, whereas cadrille relies primarily on low-level extrusion-based operations, often resulting in geometrically less faithful outputs.
These results highlight that beyond syntactic validity, the availability of expressive and semantically aligned CAD operations is crucial for faithful image-based CAD model generation.
5.6 Limitations
Although our method demonstrates considerable flexibility, it still faces challenges in generating models that require precise geometric constraints. While input text or images can explicitly specify numerical parameters (e.g., dimensions), current LLMs and VLMs exhibit limited understanding and reasoning abilities of spatial geometric relationships and topological dependencies. This often results in outputs that satisfy local constraints (e.g., correct height or diameter of a cup) but violate global consistency (e.g., the cup’s handle is inside the cup).
Due to limitations in the current code-generation models, the approach underperforms in generating CAD models involving complex mathematical computations, such as involute gears. These tasks, requiring calculations of fundamental geometric parameters, meshing relationships, or strength validation, cannot be performed well due to the complexity of generating correct spatial geometry calculation code.
6 Conclusion and Future Work
We propose CADDesigner, a novel framework for CAD modeling code generation, combined with an explicit context imperative paradigm to produce high-quality CAD modeling scripts. Experimental results demonstrate that CADDesigner achieves state-of-the-art performance in the domain of conceptual CAD design. CADDesigner is particularly effective for rapid prototyping, enabling the generation of innovative CAD models and the construction of complex, high-quality datasets that align textual descriptions with corresponding CAD modeling code.
In future work, we plan to extend CADDesigner by incorporating learning-based methods that accept point clouds and B-rep data as input, allowing the system to learn geometric and topological constraints through data-driven training. Furthermore, integrating models specialized in spatial geometric reasoning may enhance CADDesigner’s ability to compute intricate CAD parameters, such as those required for designing gears, crankshafts, and other mechanically constrained components.